# Chiral Gapless Spin Liquid in Hyperbolic Space

### v2

@@ B.2 Comparison with Exact Diagonalization and Hyperbolic Band Theory results
The phases $A_{x}$, $A_{y},$ and $A_{z}$ are realized when one of the couplings constants $J_{\alpha}$ (with $\alpha=x,y,z$) is larger than the sum of the other two. These phases contain the point where one of the coupling dominates, i.e., is much larger than the others. Without loss of generality, we consider $J_{z}\gg J_{x},J_{y}$, $J_{x}=J_{y}$ and $J_{\alpha}>0\;\forall\alpha$. In the limit $J_{x}=J_{y}=0$, the ground state consists of decoupled Ising dimers. If we contract these dimers to a point, the effective lattice becomes a hyperbolic $\{6,4\}$ lattice (Fig. C.1(b)). We can investigate the region around this limit using perturbation theory.
@@
    Each link contained in the path gives a contribution (i) $2J_{x}u_{jk}$ if $(j,k)$ is a weak link, (ii) $2J_{z}u_{jk}/(\omega^{2}+4J_{z}^{2})$ if $(j,k)$ is a strong link and (iii) $\omega/(\omega^{2}+4J_{z}^{2})$ if $(j,k)$ is a strong link that is attached to the path with only one site. The total amplitude of each path is then obtained by multiplying every contribution and integrating this product over the frequency range $\omega\in[0,\infty)$.
@@
Note that self-retracing paths give flux-independent contributions because each link $(j,k)$ is traversed once in every direction, giving a constant $u_{jk}u_{kj}=-1$. Moreover, the contributions of a path of odd length cancels out with the same path traversed in opposite direction. Thus, only non-contractible even-length paths give non-trivial contributions.
At lowest order, there are three candidate paths of length 16, two paths along two nonagons sharing a weak $J_{x}$- or $J_{y}$-link, and one along the boundary of two nonagons sharing a strong link, respectively, shown in Fig. C.1(a). Their contributions are
$I_{\text{weak}}$ $=\int_{0}^{\infty}\frac{d\omega}{2\pi}(2J_{x})^{4}(2J_{y})^{4}((2J_{x})^{2}+(2J_{y})^{2})\left(\frac{2J_{z}}{\omega^{2}+4J_{z}^{2}}\right)^{6}\left(\frac{\omega}{\omega^{2}+4J_{z}^{2}}\right)^{4}W_{9}W_{9^{\prime}}$
$=-\frac{143J_{x}^{4}J_{y}^{4}(J_{x}^{2}+J_{y}^{2})}{2^{17}J_{z}^{9}}W_{9}W_{9^{\prime}}$ (C.1)
@@
$I_{\text{strong}}$ $=\int_{0}^{\infty}\frac{d\omega}{2\pi}(2J_{x})^{6}(2J_{y})^{6}\left(\frac{2J_{z}}{\omega^{2}+4J_{z}^{2}}\right)^{4}\left(\frac{\omega}{\omega^{2}+4J_{z}^{2}}\right)^{8}W_{9}W_{9^{\prime\prime}}$
$=-\frac{91J_{x}^{6}J_{y}^{6}}{2^{20}J_{z}^{11}}W_{9}W_{9^{\prime\prime}}.$ (C.2)
The first contribution is at 10th order, the second at 12th order. The sum of these terms over the whole lattice yields the 12th order perturbative Hamiltonian (5) of the Main Text. This result is consistent with our observation that the creation of a pair of vortices costs a finite amount of energy: Flipping any link in the vortex-free or vortex-filled sector increases total energy since $W_{9}W_{9^{\prime}}=(\pm\mathrm{i})(\pm\mathrm{i})\to(\mp\mathrm{i})(\mp\mathrm{i})=-1$ yields a positive contribution when inserted into the effective Hamiltonian (5). The resulting model is a hyperbolic version of the Kitaev Toric Code [35].
@@
In this section we provide further details on the local Chern marker (LCM) [7], which we used in order to study the chiral character of the gapless phase. Let |$\ket{\psi_{n}}$ denote the eigenstates corresponding to energy $E_{n}$ of the Majorana Hamiltonian in the vortex-free configuration. Furthermore, let $\hat{P}$ be a projector onto the states with energy below an energy cutoff $\mu$,
@@
Then, the local Chern marker is defined as
@@
Figure D.1: (a) The three regions $A,B$, and $C$ in counterclockwise order we used to compute the local Chern marker. (b) The dependence of the LCM on the radius up to which sites are taken into account and the energy cutoff $\mu$ at the isotropic point (top) and in the gapped phase (bottom). The LCM is symmetric with respect to the axis $\mu/\mu_{\text{max}}=0.5$. The arrow indicates the parameters we used for computing the LCM phase diagram in Fig. 4.
In our calculation, the value of the Chern number is not quantized to integer values. While the Chern number is expected to be quantized if the Fermi level results in a band gap between two topological bands, the LCM can possibly take other values when there is no band gap as in the present case. Furthermore, the LCM calculation can be affected by finite size and edge effects in the hyperbolic plane.
@@
In order to further investigate the chiral nature of the gapless phase, we consider the time evolution for the Majorana Hamiltonian in the vortex-free and vortex-filled configuration at the isotropic point $J_{x}=J_{y}=J_{z}=1/3$. For this, we create a Gaussian wave packet with $\braket{H}=0$ centered at the intersection of the positive $x$-axis with the boundary of the sample,
@@
Figure E.1: (a) Initialized wave packet at $t=0$ localized around the rightmost boundary point of the sample in the gapless chiral phase at the isotropic point in the vortex-free sector. (b)-(e) Propagation of a wave packet using the time evolution operator $e^{-\mathrm{i}Ht}$ and time steps $dt=2000$. The center of the packet clearly propagates counterclockwise along the boundary. (f) Same wave packet as in (a) in the vortex-filled sector. (g)-(j) The packet propagates in the opposite direction with the same angular velocity as in the flux-free case because both configurations are related via time-reversal symmetry.
The time evolution is given by the operator $\hat{U}(t)=\exp[-\mathrm{i}\hat{H}t]$, where $\hat{H}$ is the free Majorana Hamiltonian in the vortex-free or vortex-filled configuration. The time evolution of the state |$\ket{\Psi_{\text{init}}}$ on a droplet with 21754 sites is shown in Fig. E.1(a)-(e) for the vortex-free sector. We observe that the center of the packet propagates along the boundary counterclockwise. Moreover, we see that the wave packet disperses and leaks into the bulk. Since the eigenspectrum of the Hamiltonian is gapless in the chiral phase, the wave packet excitation does not necessarily consist solely of states which are localized at the boundary. Thus, such behavior is expected. In the flux-filled sector the packet propagates in opposite direction (Fig. E.1(f)-(j)), as these configurations are related via time reversal.

### v4

@@ C.1 Derivation of the effective Hamiltonian for vison excitations in the dimer limit
The phases $A_{x}$, $A_{y},$ and $A_{z}$ are realized when one of the couplings constants $J_{\alpha}$ (with $\alpha=x,y,z$) is larger than the sum of the other two. These phases contain the point where one of the couplings dominates. Without loss of generality, we consider $J_{z}\gg J_{x},J_{y}$ and $J_{\alpha}>0\;\forall\alpha$. In the limit $J_{x}=J_{y}=0$, the ground state consists of decoupled Ising dimers. If we contract these dimers to a point, the effective lattice becomes a hyperbolic $\{6,4\}$ lattice (Fig. C.1(b)). We can investigate the regime around this limit using perturbation theory.
@@
    Each link contained in the path gives a contribution (i) $2J_{x}u_{jk}$ if $(j,k)$ is a weak link, (ii) $2J_{z}u_{jk}/(\omega^{2}+4J_{z}^{2})$ if $(j,k)$ is a strong link and (iii) $\omega/(\omega^{2}+4J_{z}^{2})$ if $(j,k)$ is a strong link that is attached to the path with only one site. The total amplitude of each path is then obtained by multiplying every contribution and integrating this product over the frequency $\omega$.
@@
Note that self-retracing paths give flux-independent contributions because each link $(j,k)$ is traversed once in every direction, giving a constant $u_{jk}u_{kj}=-1$. Moreover, the contributions of a path of odd length cancels out with the same path traversed in opposite direction. Thus, only noncontractible even-length paths give nontrivial contributions.
There are three candidate paths of length 16, two paths along two nonagons sharing a weak $J_{x}$ or $J_{y}$ link, and one along the boundary of two nonagons sharing a strong link, respectively, shown in Fig. C.1(a). Their contributions are
$I_{\text{weak},x}$ $=2\cdot\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\frac{1}{2}\,(2J_{x})^{4}(2J_{y})^{6}\left(\frac{2J_{z}}{\omega^{2}+4J_{z}^{2}}\right)^{6}\left(\frac{\omega}{\omega^{2}+4J_{z}^{2}}\right)^{4}W_{p}W_{p^{\prime}}=\frac{143J_{x}^{4}J_{y}^{6}}{2^{16}J_{z}^{9}}W_{p}W_{p^{\prime}},$ (C.1a)
$I_{\text{weak},y}$ $=2\cdot\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\frac{1}{2}\,(2J_{x})^{6}(2J_{y})^{4}\left(\frac{2J_{z}}{\omega^{2}+4J_{z}^{2}}\right)^{6}\left(\frac{\omega}{\omega^{2}+4J_{z}^{2}}\right)^{4}W_{p}W_{p^{\prime}}=\frac{143J_{x}^{6}J_{y}^{4}}{2^{16}J_{z}^{9}}W_{p}W_{p^{\prime}}$ (C.1b)
@@
$I_{\text{strong},z}$ $=2\cdot\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\frac{1}{2}\,(2J_{x})^{6}(2J_{y})^{6}\left(\frac{2J_{z}}{\omega^{2}+4J_{z}^{2}}\right)^{4}\left(\frac{\omega}{\omega^{2}+4J_{z}^{2}}\right)^{8}W_{p}W_{p^{\prime\prime}}=\frac{91J_{x}^{6}J_{y}^{6}}{2^{19}J_{z}^{11}}W_{p}W_{p^{\prime\prime}}.$ (C.1c)
Note that the factors of $2$ in front of the integrals come from tracing the paths in forward and backward direction. The $x$ and $y$ contributions are at 10th order, the $z$ contribution is at 12th order. Since these contributions only represent the respective leading order, the contributions by the strong links can be discarded to obtain the 10th order effective Hamiltonian. One finds
$\hat{H}_{\text{eff}}^{(10)}=\text{const}+\sum_{\langle p,p^{\prime}\rangle_{\mathrm{weak},x}}\frac{143\,J_{x}^{4}\,J_{y}^{6}}{2^{16}\,J_{z}^{9}}\,\hat{W}_{p}\hat{W}_{p^{\prime}}+\sum_{\langle p,p^{\prime}\rangle_{\mathrm{weak},y}}\frac{143\,J_{x}^{6}\,J_{y}^{4}}{2^{16}\,J_{z}^{9}}\,\hat{W}_{p}\hat{W}_{p^{\prime}},$ (C.2)
where we sum over two neighboring plaquettes $p$ and $p^{\prime}$, adjacent to either a weak $x$-type or $y$-type link.
Regarding the dimer limit, this result confirms our conjecture that the ground state lies within the uniform flux sector: If $W_{p}=\pm\mathrm{i}$ for all plaquettes $p$, the contribution of two neighboring plaquettes acquires the sign (±$(\pm\mathrm{i})(\pm\mathrm{i})=-1$. Thus, the perturbative Hamiltonian is minimized for the uniform flux sector. The resulting model is a hyperbolic version of the Kitaev Toric Code [40].
C.2 Comparison with numerical results and estimatation of finite-size effects
Refer to caption (a)(b)(c)
Figure C.2: (a) OBC cluster with a vison pair created by flipping the $z$ link (blue) between the $W_{p}=-\mathrm{i}$ plaquettes. The vison pair is surrounded by 6 $x$-type links, 6 $y$-type links, as well as 4 $z$-type links. (b)/(c) Comparison of the energy of a vison pair computed by diagonalizing the Majorana Hamiltonian for different system size in the limit $J_{x}=J_{y}=j\ll J_{z}$ with the analytical perturbation theory result of the dimer limit. Panel (b) shows the data normalized to $(j/J_{z})^{10}$ while (c) shows the plain data on a logarithmic scale. The numerically obtained energy differences converge towards the analytical result for $j/J_{z}\rightarrow 0$. Finite-size effects increase with larger $j$.
We compare this analytical result to numerical exact diagonalization (ED) calculations for finite-size droplets of different sizes. Starting from the uniform $\pi/2$ flux ground state, we create a pair of visons by flipping a strong link in the center of the droplet. This leads to two plaquettes with $W_{p}=\mathrm{i}$ surrounded by plaquettes $W_{p}=-\mathrm{i}$ (compare Fig. C.2 (a)).
Due to the vison pair, the energy of the state is increased with respect to the ground state. If we place ourselves in the dimer limit, i.e. $J_{x}=J_{y}=j\ll J_{z}$, we can compute the energy difference $\Delta E_{\mathrm{lim}}$ to the ground state analytically using equation (C.2). There is an energy contribution for each pair of neighboring plaquettes with opposite flux if the plaquettes are connected by a weak link. Plaquette pairs with the same flux do not contribute, and the corrections from opposite flux plaquettes connected by a strong link are negligible. Therefore, we have to consider $12$ contributions in total (6 from $x$ link boundaries to $+\mathrm{i}$ plaquettes, 6 from $y$ link boundaries to $+\mathrm{i}$ plaquettes), leading to an energy increase of
$\Delta E_{\mathrm{lim}}=2\cdot 12\cdot\frac{143\,J_{z}}{2^{16}}\left(\frac{j}{J_{z}}\right)^{10}\approx$0.052\,368$\,J_{z}\,\left(\frac{j}{J_{z}}\right)^{10}\qquad\text{for }\frac{j}{J_{z}}\rightarrow 0$ (C.3)
compared to the ground state.
We check this result numerically by constructing and diagonalizing the Majorana Hamiltonian for this flux configuration. Summing the eigenvalues of the Majorana Hamiltonian in the prescribed way yields the eigenenergy of the many-body state. We compute this energy for the ground state as well as for the two-vison state for different values of $\frac{j}{J_{z}}$. Note that this computation requires caution as the energy difference between both states scales with $(\frac{j}{J_{z}})^{10}$, which necessitates high computational precision. For example, for $\frac{j}{J_{z}}=$0.1$$, the numerical precision of the eigenvalues must be at least $10^{-13}$ or $10^{-14}$. The energy difference $\Delta E$ resulting from these computations is compared with the analytical result $\Delta E_{\mathrm{lim}}$ in figure C.2 (b) and (c). Panel (b) shows the result normalized to $(\frac{j}{J_{z}})^{10}$, while panel (c) depicts the actual energy increase on a logarithmic scale. As we can see, the numerical results approach the analytical value as $\frac{j}{J_{z}}\rightarrow 0$, which indeed confirms that the perturbation theory for the vison excitations in the dimer limit holds.
The numerical data were obtained for different sizes of the OBC cluster (c.f. Fig. C.2 (b), (c)), which enables us to gauge the finite-size effect in this type of calculations. As it turns out, the finite-size effects on excitation energies impact the result significantly for larger values of $\frac{j}{J_{z}}$. This means especially, that computations at the isotropic point ($j/J_{z}=1$) must be assumed to suffer from certain finite-size effects for the available system sizes. This holds certainly also for the quantitative findings of figure A.1.
@@ C.2 Comparison with numerical results and estimatation of finite-size effects
In this section we provide further details on the local Chern marker (LCM) [7], which we used in order to study the chiral character of the gapless phase. Let |$\ket{\psi_{n}}$ denote the eigenstate corresponding to energy $E_{n}$ of the Majorana Hamiltonian in the $-\frac{\pi}{2}$ flux configuration. Furthermore, let $\hat{P}$ be a projector onto all eigenstates with energy below an energy cutoff $\mu$,
@@
Then, the LCM is defined as
@@
Figure D.1: (a) The three regions $A,B$, and $C$ in counterclockwise order we used to compute the local Chern marker. (b) The dependence of the LCM on the radius up to which sites are taken into account and the energy cutoff $\mu$ at the isotropic point (top, $J_{x}=J_{y}=J_{z}=1/3$) and in the gapped phase (bottom, $J_{x}=J_{y}=1/5$, $J_{z}=3/5$). The LCM is symmetric with respect to the axis $\mu/\mu_{\text{max}}=0.5$. The arrow indicates the parameters we used for computing the LCM phase diagram in Fig. 4.
In our calculation, the value of the LCM is not quantized to integer values. While the Chern number is expected to be quantized if the Fermi level results in a band gap between two topological bands, the LCM as a computational tool can possibly take other values, especially in the absence of a band gap as in the present case. Furthermore, the LCM calculation can be affected by finite-size and edge effects in the hyperbolic plane.
@@
In order to further investigate the chiral nature of the gapless phase, we consider the time evolution for the Majorana Hamiltonian in the ±$\pm\frac{\pi}{2}$ flux configuration at the isotropic point $J_{x}=J_{y}=J_{z}=1/3$. For this, we create a Gaussian wave packet with $\braket{H}=0$ centered at the intersection of the positive $x$ axis with the boundary of the sample,
@@
Figure E.1: (a) Initialized wave packet at $t=0$ localized around the rightmost boundary point of the sample in the gapless chiral phase at the isotropic point in the $-\frac{\pi}{2}$ flux sector. (b)-(e) Propagation of a wave packet using the time evolution operator $e^{-\mathrm{i}\hat{H}t}$ and time steps $dt=800$. The center of the packet clearly propagates counterclockwise along the boundary. (f) Same wave packet as in (a) in the $+\frac{\pi}{2}$ flux sector. (g)-(j) The packet propagates in the opposite direction with the same angular velocity as in the $-\frac{\pi}{2}$ flux case because both configurations are related via time-reversal symmetry.
The time evolution is given by the operator $\exp[-\mathrm{i}\hat{H}t]$, where $\hat{H}$ is the free Majorana Hamiltonian in the ±$\pm\frac{\pi}{2}$ flux configuration. The time evolution of the state |$\ket{\Psi_{\text{init}}}$ on a droplet with 21754 sites is shown in Fig. E.1(a)-(e) for the $-\frac{\pi}{2}$ flux sector. We observe that the center of the packet propagates along the boundary counterclockwise. Moreover, we see that the wave packet disperses and leaks into the bulk. Since the eigenspectrum of the Hamiltonian is gapless in the chiral phase, the wave packet excitation does not necessarily consist solely of states which are localized at the boundary. Thus, such behavior is expected. In the $+\frac{\pi}{2}$ flux sector the packet propagates in opposite direction (Fig. E.1(f)-(j)), as these configurations are related via time reversal.